# Chiral Heisenberg Gross-Neveu-Yukawa criticality: honeycomb vs. SLAC fermions

### v1

@@ III.2.1 Weak coupling Dirac semi-metal
We observe a stable DSM phase up to the critical point, beyond which AFM order emerges and the fermions acquire mass. The emergence of long-range AFM order can be tracked by measuring the spin structure factor ${S_{\text{AFM}}(\mathbf{k})\equiv\sum_{\mathbf{r}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\langle\mathbf{S}(\mathbf{r})\cdot\mathbf{S}(\mathbf{0})\rangle/L^{2}}$, where the spin ${\mathbf{S}_{\mathbf{r}}=\frac{1}{2}a_{\mathbf{r}\alpha}^{\dagger}\bm{\sigma}_{\alpha\beta}a_{\mathbf{r}\beta}}$ at position $\mathbf{r}$, orbital $a$, and $\bm{\sigma}$ denotes the vector of the three Pauli matrices. The squared magnetization then is obtained from ${m^{2}=S_{\text{AFM}}(\mathbf{k}=0)/L^{2}}$. Figure 3 shows the evolution of the finite size order parameter as a function of the Coulomb repulsion, which indicates the quantum phase transition to be located at ${U_{c}\approx 6}$. The inset shows the finite size extrapolation of the single-particle gap $\Delta_{\text{sp}}(\mathbf{k})$ obtained from a fit to the asymptotic long imaginary-time behavior of the single particle Green’s function ${G(\mathbf{k},\tau)=\langle a_{\mathbf{k}}^{\dagger}(\tau)\,a_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\text{sp}}(\mathbf{k})]}$. The gap vanishes faster than ${1/L}$ consistent with simulations of local models on a torus [27, 57, 58] below the critical coupling ${U\leq U_{c}}$, such that the Dirac fermions remain massless in the TDL and the gap opens with the onset of order. We detect no anisotropy in the single particle and spin Green’s functions, all supporting a stable semi-metal fixed point. Potential long-range spin fluctuations as a consequence of dynamically induced super-exchange mediated by the long range-hopping apparently do not destabilize the DSM.
@@
The locality of the Dirac operator is recovered in the infinite volume limit for most of the BZ, but the singular boundary [59]. We thus monitor the dominant single particle channel and AFM correlations for interfering low-energy modes induced by interactions, in particular close to the boundary of the BZ. Such excitations (doublers) could potentially alter the degeneracy of the fermions at the Fermi level and consequently the universality class of the quantum phase transition. The single-particle spectral function ${A(\mathbf{k},\omega)=-{\rm Im}\,G(\mathbf{k},\omega)/\pi}$ close to the critical point as obtained from stochastic analytical continuation [60, 61] is presented in Fig. 4 for momenta along the path in the BZ sketched in Fig. 2(a). The dispersion of the non-interacting system is indicated by the white dashed line. The excitation gap at $\Gamma$ is the one most affected by finite size corrections, yet scales predominately as ${\Delta_{\text{sp}}\propto L^{-z}}$. The finite size corrections of the single-particle excitations away from vanishing momentum are significantly smaller leading to a scaling faster than $L^{-1}$ [62, 58]. Note that the quasiparticle weight $Z$ (the intensity in Fig. 4) actually vanishes at ${\mathbf{k}\to\Gamma}$ at the critical point as ${Z\sim(U-U_{c})^{-z\nu}}$ in the TDL [27, 63]. The discontinuous dispersion at the BZ boundary appears to be smoothed upon the introduction of interactions and the single particle excitations close to the boundary of the BZ converge to a finite values at rather high energies. We indicate the finite size extrapolated values of the excitations at the momenta ${M^{\prime}=(\pi-2\pi/L,0)}$ and ${X^{\prime}=(\pi-2\pi/L,\pi-2\pi/L)}$ closest to the Brillouin zone boundary. No additional zero modes introduced by correlations as reflected in the interacting single-particle spectrum, nor the spin spectrum as discussed below at criticality.
@@
The envelope of the dispersion close to criticality suggests there is only weak renormalization of the Fermi velocity. This is in accordance with results from appropriately chosen estimators on the honeycomb and other lattices, as well as for ${N=4}$ [62, 16, 27]. In the absence of the knowledge of the level structure from conformal field theory we cannot determine the true Fermi velocity at the critical point [64, 65, 58]. We can however track the Fermi velocity approaching the critical point from the noninteracting DSM fixed point, where the speed of light $v_{F}$ is known, as shown in Fig. 5.
@@
Here we extract the fermionic excitations velocity on the torus from fits of the relativistic energy momentum relation to the envelope of the excitations gaps $\Delta(\mathbf{k})$ in the range ${k<1.3}$, such that for different system sizes
@@ III.2.2 Strong coupling AFM
At strong coupling ${U\gg v_{F}}$ the system spontaneously orders antiferromagnetically, driven by an effective super-exchange coupling, which results from predominantly second order hopping processes as ${J\sim t^{2}/U\sim 1/Ur^{2}}$. As the fermions are gapped out we can approximate the mostly localized ground state as Heisenberg ${S=1/2}$ antiferromagnetic via the Holstein-Primakoff transformation to linear order [66]. The resulting bosonic Hamiltonian in the basis ${(a_{\mathbf{k}}^{\dagger},b_{\mathbf{k}}^{\phantom{\dagger}})}$, where the two orbitals are associated with the sublattices at momentum $\mathbf{k}$, has the form
@@
where we have used the explicit finite size expression for ${t(r)}$ from Eq. (3). The Hamiltonian (12) may be diagonalized by a Bogoliubov transformation, which yields the eigenvalues
@@
=$\stackrel{{\scriptstyle L\to\infty}}{{=}}4\sum_{r=1}^{\infty}\frac{1}{r^{2}}=4\zeta(2)=\frac{2\pi}{3}\;.$ (15)
@@
whereas for finite size systems we use ${\varepsilon_{\text{LSWT}}(\mathbf{k})=\sqrt{J_{\mathbf{0}}^{2}-J_{\mathbf{k}}^{2}}}$. While retaining all lattice symmetries, the resulting dispersion is anisotropic as a consequence of the major axis coupling and shown in the inset of Fig. 6. In contrast, in fully connected long range models the magnon dispersion is isotropic. For small momenta the dispersion follows $\sqrt{k}$ similar to fully connected long range models [67], but for the major axis anisotropy. At strong coupling Eq. (16) approximates the true spin wave dispersion. This is shown for ${U=20}$, ${N=4}$ along the cuts ${(k,0)}$ and ${(k,k)}$ in Fig. 7, where we extracted the spin gap $\Delta_{\sigma}$ obtained from a fit to the asymptotic long imaginary-time behavior of the unequal-time spin-spin correlation function ${S(\mathbf{k},\tau)=\langle S_{\mathbf{k}}^{\dagger}(\tau)\,S_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\sigma}(\mathbf{k})]}$.
@@
where $v_{\phi}$ denotes the spin-wave velocity fitted to the data. This unwarping enables us to remove the leading non-linear behavior from the finite size data, yet retains the correct limit for vanishing momentum [58]. The then approximately linear dispersion allows for the extrapolation of the spin gaps to zero, which is consistent with the emergence of gapless Goldstone modes. While anomalous in their $\sqrt{k}$-dispersion, compared to a fully connected effective long-range Heisenberg model, the behavior also implies that the dynamically generated super-exchange coupling decays significantly faster than would be required for a gapped Goldstone mode [67] which has been claimed for the ${N=4}$ SLAC-Hubbard model in [52].
@@
Figure 7: The momentum dependence of the QMC spin gap $\Delta_{\sigma}(k)$ at strong coupling ${U=20}$ along the paths $(0,0)$ to ${(\pi,0)}$ and ${(\pi,\pi)}$ is well reproduced by the finite size LSWT (orange ${L=19}$, gray ${L\to\infty}$).
Figure 8: Finite size extrapolation the unwarped QMC data at strong coupling ${U=20}$ (c.f. inset) via the LSWT spin-wave dispersion ${\varepsilon_{\text{LSWT}}}$ to remove the leading nonlinear behavior. The linear finite size extrapolation of the smallest unwarped spin gaps ${\hat{\Delta}_{\sigma}(k\to 0)}$ to zero provides evidence for gapless, anomalous Goldstone modes.
@@ III.2.3 Criticality
The question remains whether the effective long-range coupling alters the properties at the critical point which then would no longer be representative of the ${N=8}$ chiral Heisenberg universality class. First, we observe finite size single particle and spin gaps that vanish in the TDL (c.f. inset in Fig. 3, Fig. 8 and Fig. 9) consistent with gapless excitations in the two relevant sectors. Across criticality we further expect the fermionic Dirac spectrum to be replaced by its bosonic counterpart, such that the primary excitations share the same velocity. In Fig. 9 we can see that at criticality and in contrast to Fig. 7 at strong coupling, the spin wave dispersion is isotropic, with an excitation velocity approaching the non-interacting Fermi velocity as indicated by the gray dashed line. This also provides an independent indicator that there is little to no renormalization of the Fermi velocity. Where the finite size dependence of the spin gaps at zero momentum is clear, for finite momenta this is less the case. Whether the shift to higher energies off the dashed line indicating the non-interacting Fermi velocity is a finite size effect or a persistent remnant of the dynamically induced long-range spin-exchange remains open. Again, we find no evidence for super-luminal magnon excitations. The recovered isotropy and similar velocities of the spin excitations compared to the fermionic single particle spectrum are in accordance with the expectations for the GNY-type transition.
@@ IV.1 Proximity to asymptotic scaling
where $\mathbf{b}_{1}$, $\mathbf{b}_{2}$ denote the reciprocal lattice vectors and $\mathbf{Q}$ is the momentum at which order develops across the transition [70, 71]. The correlation function $C$ must be associated with the order parameter of the phase transition. Here, the correlation function corresponds to the spin structure factor ${S_{\text{AFM}}(\mathbf{Q})/L^{2}=m^{2}}$ and we will denote the CR as $R_{m^{2}}$. The critical point can be inferred from the intersections of different pairs of system sizes where ${R_{m^{2}}(u,L)=R_{m^{2}}(u,aL+b)}$. Here ${a,b\in\mathbb{N}}$ with a typical choice of ${a=2}$ in order to express the dominant scaling behavior and ${b=0,1}$ to account for commensurate lattice sizes. In the TDL at the critical coupling the CR approaches ${\lim_{L\to\infty}R_{m^{2}}(0,L)=R_{m^{2}}^{*}}$ [72, 73].
@@ IV.2 Finite size scaling analysis
$m^{2}(R_{m^{2}},L)=L^{1-\eta_{\phi}}f_{0}^{m}(R_{m^{2}})\;,$ (20)
where we assume ${z=1}$ [74, 11]. Fits to the data are shown in Fig. 14, where the insets illustrate the decay ${m^{2}\sim L^{1-\eta_{\phi}}}$ using the extracted $\eta_{\phi}$ from the main panels.
@@
again under the assumption of ${z=1}$.
@@ V Conclusion
[…]
@@ Appendix A Hopping amplitude of SLAC Fermions
Here we calculate the discrete Fourier transformation of the hopping matrix elements for SLAC fermions on lattices of even and odd linear dimension $L$ which include the Dirac point ${k=0}$. Let us compute real-space representation of the linear dispersion ${\varepsilon(k)=k}$ in one dimension
@@
where ${k=2\pi m/L}$ and ${m=-M,\ldots,M\in\mathbb{N}}$ with ${M=(L-1)/2}$ for odd $L$ system sizes. In the following we will use the exponential sum formula to replace the sum with a fraction
$\sum_{k=0}^{N-1}k\,\mathrm{e}^{\mathrm{i}kx}$ $=\frac{\mathrm{d}}{\mathrm{d}x}\sum_{k=0}^{N-1}\,\mathrm{e}^{\mathrm{i}kx}$
$=\frac{\mathrm{d}}{\mathrm{d}x}\frac{1-\mathrm{e}^{\mathrm{i}Nx}}{1-\mathrm{e}^{\mathrm{i}x}}$
$=\frac{N\,\mathrm{e}^{\mathrm{i}Nx}\left(1-\mathrm{e}^{\mathrm{i}x}\right)-\mathrm{e}^{\mathrm{i}x}\left(1-\mathrm{e}^{\mathrm{i}Nx}\right)}{\left(1-\mathrm{e}^{\mathrm{i}x}\right)^{2}}~{},$ (23)
@@
$\varepsilon(x)$ $=\frac{1}{L}\sum_{k}k\,\left(\,\mathrm{e}^{\mathrm{i}kx}-\,\mathrm{e}^{-\mathrm{i}kx}\right)$
$=\frac{2\mathrm{i}}{L}\sum_{k}k\sin\left(kx\right)$
$=\frac{2\mathrm{i}}{L}\;\mathrm{Im}\sum_{k}k\,\mathrm{e}^{\mathrm{i}kx}$
@@
$\varepsilon(x)$ $=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\frac{L+1}{2}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\frac{L+1}{2}x}\right)}{\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)}{\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\mathrm{e}^{\mathrm{i}\pi x}(-2\mathrm{i})\,\sin\left(\frac{\pi}{L}x\right)-\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}\right]$
$=\frac{4\mathrm{i}\pi}{L^{2}}(-1)^{x}\frac{\frac{L+1}{2}\,(-2)\,\sin\left(\frac{\pi}{L}x\right)-\sin\left(\frac{\pi}{L}x\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}$
$=\frac{\mathrm{i}\pi}{L}\frac{(-1)^{x}}{\sin\left(\frac{\pi}{L}x\right)}\;.$ (25)

### v2

@@ III.2.1 Weak coupling Dirac semi-metal
We observe a stable DSM phase up to the critical point, beyond which AFM order emerges and the fermions acquire mass. The emergence of long-range AFM order can be tracked by measuring the spin structure factor $S_{\text{AFM}}(\mathbf{k})\equiv\sum_{\mathbf{r}}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\langle\mathbf{S}(\mathbf{r})\cdot\mathbf{S}(\mathbf{0})\rangle/L^{2}$}, where the spin $\mathbf{S}_{\mathbf{r}}=\frac{1}{2}a_{\mathbf{r}\alpha}^{\dagger}\bm{\sigma}_{\alpha\beta}a_{\mathbf{r}\beta}$} at position $\mathbf{r}$, orbital $a$, and $\bm{\sigma}$ denotes the vector of the three Pauli matrices. The squared magnetization then is obtained from $m^{2}=S_{\text{AFM}}(\mathbf{k}=0)/L^{2}$}. Figure 3 shows the evolution of the finite size order parameter as a function of the Coulomb repulsion, which indicates the quantum phase transition to be located at $U_{c}\approx 6$}. The inset shows the finite size extrapolation of the single-particle gap $\Delta_{\text{sp}}(\mathbf{k})$ obtained from a fit to the asymptotic long imaginary-time behavior of the single particle Green’s function $G(\mathbf{k},\tau)=\langle a_{\mathbf{k}}^{\dagger}(\tau)\,a_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\text{sp}}(\mathbf{k})]$}. The gap vanishes faster than 1/L{1/L} consistent with simulations of local models on a torus [27, 59, 60] below the critical coupling $U\leq U_{c}$}, such that the Dirac fermions remain massless in the TDL and the gap opens with the onset of order. We detect no anisotropy in the single particle and spin Green’s functions, all supporting a stable semi-metal fixed point. Potential long-range spin fluctuations as a consequence of dynamically induced super-exchange mediated by the long range-hopping apparently do not destabilize the DSM.
@@
The locality of the Dirac operator is recovered in the infinite volume limit for most of the BZ, but the singular boundary [61, 62, 44]. We thus monitor the dominant single particle channel and AFM correlations for interfering low-energy modes induced by interactions, in particular close to the boundary of the BZ. Such excitations (doublers) could potentially alter the degeneracy of the fermions at the Fermi level and consequently the universality class of the quantum phase transition. The single-particle spectral function $A(\mathbf{k},\omega)=-{\rm Im}\,G(\mathbf{k},\omega)/\pi$} close to the critical point as obtained from stochastic analytical continuation [63, 64] is presented in Fig. 4 for momenta along the path in the BZ sketched in Fig. 2(a). The dispersion of the non-interacting system is indicated by the white dashed line. The excitation gap at $\Gamma$ is the one most affected by finite size corrections, yet scales predominately as $\Delta_{\text{sp}}\propto L^{-z}$}. The finite size corrections of the single-particle excitations away from vanishing momentum are significantly smaller leading to a scaling faster than $L^{-1}$ [65, 60]. Note that the quasiparticle weight $Z$ (the intensity in Fig. 4) actually vanishes at $\mathbf{k}\to\Gamma$} at the critical point as $Z\sim(U-U_{c})^{-z\nu}$} in the TDL [27, 66]. The discontinuous dispersion at the BZ boundary appears to be smoothed upon the introduction of interactions and the single particle excitations close to the boundary of the BZ converge to a finite values at rather high energies. We indicate the finite size extrapolated values of the excitations at the momenta $M^{\prime}=(\pi-2\pi/L,0)$} and $X^{\prime}=(\pi-2\pi/L,\pi-2\pi/L)$} closest to the Brillouin zone boundary. No additional zero modes introduced by correlations as reflected in the interacting single-particle spectrum, nor the spin spectrum as discussed below at criticality.
@@
The envelope of the dispersion close to criticality suggests there is only weak renormalization of the Fermi velocity. This is in accordance with results from appropriately chosen estimators on the honeycomb and other lattices, as well as for N=4{N=4} [65, 16, 27]. In the absence of the knowledge of the level structure from conformal field theory we cannot determine the true Fermi velocity at the critical point [67, 68, 60]. We can however track the Fermi velocity approaching the critical point from the noninteracting DSM fixed point, where the speed of light $v_{F}$ is known, as shown in Fig. 5.
@@
Here we extract the fermionic excitations velocity on the torus from fits of the relativistic energy momentum relation to the envelope of the excitations gaps $\Delta(\mathbf{k})$ in the range k<1.3{k<1.3}, such that for different system sizes
@@ III.2.2 Strong coupling AFM
At strong coupling $U\gg v_{F}$} the system spontaneously orders antiferromagnetically, driven by an effective super-exchange coupling, which results from predominantly second order hopping processes as $J\sim t^{2}/U\sim 1/Ur^{2}$}. As the fermions are gapped out we can approximate the mostly localized ground state as Heisenberg S=1/2{S=1/2} antiferromagnetic via the Holstein-Primakoff transformation to linear order [69]. The resulting bosonic Hamiltonian in the basis $(a_{\mathbf{k}}^{\dagger},b_{\mathbf{k}}^{\phantom{\dagger}})$}, where the two orbitals are associated with the sublattices at momentum $\mathbf{k}$, has the form
@@
where we have used the explicit finite size expression for t(r){t(r)} from Eq. (3). The Hamiltonian (12) may be diagonalized by a Bogoliubov transformation, which yields the eigenvalues
@@
$\stackrel{{\scriptstyle L\to\infty}}{{=}}4\sum_{r=1}^{\infty}\frac{1}{r^{2}}=4\zeta(2)=\frac{2\pi}{3}\;.$ (15)
@@
whereas for finite size systems we use $\varepsilon_{\text{LSWT}}(\mathbf{k})=\sqrt{J_{\mathbf{0}}^{2}-J_{\mathbf{k}}^{2}}$}. While retaining all lattice symmetries, the resulting dispersion is anisotropic as a consequence of the major axis coupling and shown in the inset of Fig. 6. In contrast, in fully connected long range models the magnon dispersion is isotropic. For small momenta the dispersion follows $\sqrt{k}$ similar to fully connected long range models [70], but for the major axis anisotropy. At strong coupling Eq. (16) approximates the true spin wave dispersion. This is shown for U=20{U=20}, N=4{N=4} along the cuts (k,0){(k,0)} and (k,k){(k,k)} in Fig. 7, where we extracted the spin gap $\Delta_{\sigma}$ obtained from a fit to the asymptotic long imaginary-time behavior of the unequal-time spin-spin correlation function $S(\mathbf{k},\tau)=\langle S_{\mathbf{k}}^{\dagger}(\tau)\,S_{\mathbf{k}}^{\phantom{\dagger}}(0)\rangle\propto\exp[-\tau\Delta_{\sigma}(\mathbf{k})]$}.
@@
where $v_{\phi}$ denotes the spin-wave velocity fitted to the data. This unwarping enables us to remove the leading non-linear behavior from the finite size data, yet retains the correct limit for vanishing momentum [60]. The then approximately linear dispersion allows for the extrapolation of the spin gaps to zero, which is consistent with the emergence of gapless Goldstone modes. While anomalous in their $\sqrt{k}$-dispersion, compared to a fully connected effective long-range Heisenberg model, the behavior also implies that the dynamically generated super-exchange coupling decays significantly faster than would be required for a gapped Goldstone mode [70] which has been claimed for the N=4{N=4} SLAC-Hubbard model in [53].
@@
Figure 7: The momentum dependence of the QMC spin gap $\Delta_{\sigma}(k)$ at strong coupling U=20{U=20} along the paths $(0,0)$ to $(\pi,0)$} and $(\pi,\pi)$} is well reproduced by the finite size LSWT (orange L=19{L=19}, gray $L\to\infty$}).
Figure 8: Finite size extrapolation the unwarped QMC data at strong coupling U=20{U=20} (c.f. inset) via the LSWT spin-wave dispersion $\varepsilon_{\text{LSWT}}$} to remove the leading nonlinear behavior. The linear finite size extrapolation of the smallest unwarped spin gaps $\hat{\Delta}_{\sigma}(k\to 0)$} to zero provides evidence for gapless, anomalous Goldstone modes.
@@ III.2.3 Criticality
The question remains whether the effective long-range coupling alters the properties at the critical point which then would no longer be representative of the N=8{N=8} chiral Heisenberg universality class. First, we observe finite size single particle and spin gaps that vanish in the TDL (c.f. inset in Fig. 3, Fig. 8 and Fig. 9) consistent with gapless excitations in the two relevant sectors. Across criticality we further expect the fermionic Dirac spectrum to be replaced by its bosonic counterpart, such that the primary excitations share the same velocity [3, 71, 58]. In Fig. 9 we can see that at criticality and in contrast to Fig. 7 at strong coupling, the spin wave dispersion is isotropic, with an excitation velocity approaching the non-interacting Fermi velocity as indicated by the gray dashed line. This also provides an independent indicator that there is little to no renormalization of the Fermi velocity. Where the finite size dependence of the spin gaps at zero momentum is clear, for finite momenta this is less the case. Whether the shift to higher energies off the dashed line indicating the non-interacting Fermi velocity is a finite size effect or a persistent remnant of the dynamically induced long-range spin-exchange remains open. Again, we find no evidence for super-luminal magnon excitations. The recovered isotropy and similar velocities of the spin excitations compared to the fermionic single particle spectrum are in accordance with the expectations for the GNY-type transition.
@@ IV.1 Proximity to asymptotic scaling
where $\mathbf{b}_{1}$, $\mathbf{b}_{2}$ denote the reciprocal lattice vectors and $\mathbf{Q}$ is the momentum at which order develops across the transition [74, 75]. The correlation function $C$ must be associated with the order parameter of the phase transition. Here, the correlation function corresponds to the spin structure factor $S_{\text{AFM}}(\mathbf{Q})/L^{2}=m^{2}$} and we will denote the CR as $R_{m^{2}}$. The critical point can be inferred from the intersections of different pairs of system sizes where $R_{m^{2}}(u,L)=R_{m^{2}}(u,aL+b)$}. Here $a,b\in\mathbb{N}$} with a typical choice of a=2{a=2} in order to express the dominant scaling behavior and b=0,1{b=0,1} to account for commensurate lattice sizes. In the TDL at the critical coupling the CR approaches $\lim_{L\to\infty}R_{m^{2}}(0,L)=R_{m^{2}}^{*}$} [76, 77].
@@ IV.2 Finite size scaling analysis
$m^{2}(R_{m^{2}},L)=L^{-(1+\eta_{\phi})}f_{0}^{m}(R_{m^{2}})\;,$ (20)
where we assume z=1{z=1} [78, 11]. Fits to the data are shown in Fig. 14, where the insets illustrate the decay $m^{2}\sim L^{-(1+\eta_{\phi})}$} using the extracted $\eta_{\phi}$ from the main panels.
@@
again under the assumption of z=1{z=1}.
@@ V Conclusion
The system at the critical point is at least non-perturbatively renormalizable and the extracted critical exponents presented here and now in several cases of simulations with a single Dirac cone in 2+1dd appear to be in accordance with existing estimates from local models, in particular in cases where we have exact results to compare with [33, 5, 27, 28, 24, 82]. Most importantly, they significantly reduce finite size effects, which are a major limiting factor in QMC simulations. Moreover, the SLAC formulation may be modified for a variety of low energy dispersions, which might be realized in local models, but where the physics in interacting systems is hidden at small energy scales which cannot be resolved in accessible lattice sizes of local Hamiltonians.
Pathological effects of SLAC fermions in 2+1dd at T=0{T=0} cannot not be ruled out completely and more investigations are certainly called for. Yet, neither the anomalous Goldstone modes, nor the finite temperature AFM order imply that the nature of the QCP, i.e., its critical properties, differ from the chiral Heisenberg GNY universality class or the QCP found in local models. If one considers the anisotropic AFM a different type of fixed point than the AFM on the honeycomb, albeit with the same remaining symmetries, they can still share the same critical properties: The working assumption of GNY universality remains that critical properties depend on the dimensionality, the symmetry of the order parameter and number of fermions. Furthermore, long-range hopping in the Hamiltonian does not exclude a conformal fixed point. On the contrary, effective actions obtained from integrating out, e.g., gauge fields, often result in non-local, yet conformal field theories.
As such, the use of SLAC fermions in 2+1dd remains an intriguing, non-trivial issue with strong positive evidence in favor of the SLAC formulation. We attribute this to the optimal kinetic term, which produces a perfect Dirac cone across the whole BZ, rather than only for a limited portion of available momenta and as such minimizes non-relativistic contributions from the noninteracting dispersion.
@@ Appendix A Hopping amplitude of SLAC Fermions
Here we calculate the discrete Fourier transformation of the hopping matrix elements for SLAC fermions on lattices of even and odd linear dimension $L$ which include the Dirac point k=0{k=0}. Let us compute real-space representation of the linear dispersion $\varepsilon(k)=k$} in one dimension
@@
where $k=2\pi m/L$} and $m=-M,\ldots,M\in\mathbb{N}$} with M=(L1)/2{M=(L-1)/2} for odd $L$ system sizes. In the following we will use the exponential sum formula to replace the sum with a fraction
$\sum_{k=0}^{N-1}k\,\mathrm{e}^{\mathrm{i}kx}$ $=\frac{1}{\mathrm{i}}\frac{\mathrm{d}}{\mathrm{d}x}\sum_{k=0}^{N-1}\,\mathrm{e}^{\mathrm{i}kx}$
$=\frac{1}{\mathrm{i}}\frac{\mathrm{d}}{\mathrm{d}x}\frac{1-\mathrm{e}^{\mathrm{i}Nx}}{1-\mathrm{e}^{\mathrm{i}x}}$
$=-\frac{N\,\mathrm{e}^{\mathrm{i}Nx}\left(1-\mathrm{e}^{\mathrm{i}x}\right)-\mathrm{e}^{\mathrm{i}x}\left(1-\mathrm{e}^{\mathrm{i}Nx}\right)}{\left(1-\mathrm{e}^{\mathrm{i}x}\right)^{2}}~,$ (23)
@@
$\varepsilon(x)$ $=\frac{1}{L}\sum_{k\geq 0}k\,\left(\,\mathrm{e}^{\mathrm{i}kx}-\,\mathrm{e}^{-\mathrm{i}kx}\right)$
$=\frac{2\mathrm{i}}{L}\sum_{k\geq 0}k\sin\left(kx\right)$
$=\frac{2\mathrm{i}}{L}\;\mathrm{Im}\sum_{k\geq 0}k\,\mathrm{e}^{\mathrm{i}kx}$
@@
$\varepsilon(x)$ $=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\frac{L+1}{2}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}\frac{L+1}{2}x}\right)}{\left(1-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)-\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)}{\,\mathrm{e}^{\mathrm{i}\frac{2\pi}{L}x}\left(\mathrm{e}^{-\mathrm{i}\frac{\pi}{L}x}-\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}\;\mathrm{Im}\left[\frac{\frac{L+1}{2}\mathrm{e}^{\mathrm{i}\pi x}(-2\mathrm{i})\,\sin\left(\frac{\pi}{L}x\right)-\left(1-\mathrm{e}^{\mathrm{i}\pi x}\,\mathrm{e}^{\mathrm{i}\frac{\pi}{L}x}\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}\right]$
$=-\frac{4\mathrm{i}\pi}{L^{2}}(-1)^{x}\frac{\frac{L+1}{2}\,(-2)\,\sin\left(\frac{\pi}{L}x\right)-\sin\left(\frac{\pi}{L}x\right)}{-4\,\sin\left(\frac{\pi}{L}x\right)^{2}}$
$=-\frac{\mathrm{i}\pi}{L}\frac{(-1)^{x}}{\sin\left(\frac{\pi}{L}x\right)}\;.$ (25)